# Quantum fluctuation energies over a spatially inhomogeneous field background in a chiral soliton model

### v3

@@ Appendix A Calculations of Born subtractions of phase shifts by the first order differential equations
    $\left\{\begin{aligned} &v_{1}^{(1)(1,0)\prime}+v_{1}^{(1)(1,0)}\frac{h_{G+1}^{\prime}}{h_{G+1}}+\frac{G+2}{r}v_{1}^{(1)(1,0)}-\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G+1}}u_{1}^{(1)(1,0)}+\beta\sigma_{s}=0;\\ &v_{2}^{(1)(1,0)\prime}+v_{2}^{(1)(1,0)}\frac{h_{G-1}^{\prime}}{h_{G-1}}-\frac{G-1}{r}v_{2}^{(1)(1,0)}+\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G-1}}u_{2}^{(1)(1,0)}=0\text{; }\\ &u_{1}^{(1)(1,0)\prime}+u_{1}^{(1)(1,0)}\frac{h_{G}^{\prime}}{h_{G}}-\frac{G}{r}u_{1}^{(1)(1,0)}+\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G+1}}{\mathrm{~h}_{G}}v_{1}^{(1)(1,0)}+\sigma_{s}=0;\\ &u_{2}^{(1)(1,0)\prime}+u_{2}^{(1)(1,0)}\frac{h_{G}^{\prime}}{h_{G}}+\frac{G+1}{r}u_{2}^{(1)(1,0)}-\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G-1}}{\mathrm{~h}_{G}}v_{2}^{(1)(1,0)}=0;\end{aligned}\right$. (77)
    $\left\{\begin{aligned} &v_{1}^{(2)(1,0)\prime}+v_{1}^{(2)(1,0)}\frac{h_{G+1}^{\prime}}{h_{G+1}}+\frac{G+2}{r}v_{1}^{(2)(1,0)}-\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G+1}}u_{1}^{(2)(1,0)}=0;\\ &v_{2}^{(2)(1,0)\prime}+v_{2}^{(2)(1,0)}\frac{h_{G-1}^{\prime}}{h_{G-1}}-\frac{G-1}{r}v_{2}^{(2)(1,0)}+\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G-1}}u_{2}^{(2)(1,0)}-\beta\sigma_{s}=0;\\ &u_{1}^{(2)(1,0)\prime}+u_{1}^{(2)(1,0)}\frac{h_{G}^{\prime}}{h_{G}}-\frac{G}{r}u_{1}^{(2)(1,0)}+\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G+1}}{\mathrm{~h}_{G}}v_{1}^{(2)(1,0)}=0;\\ &u_{2}^{(2)(1,0)\prime}+u_{2}^{(2)(1,0)}\frac{h_{G}^{\prime}}{h_{G}}+\frac{G+1}{r}u_{2}^{(2)(1,0)}-\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G-1}}{\mathrm{~h}_{G}}v_{2}^{(2)(1,0)}-\sigma_{s}=0;\end{aligned}\right$. (78)
@@
    $\left\{\begin{aligned} &v_{1}^{(1)(2,0)\prime}+v_{1}^{(1)(2,0)}\frac{h_{G+1}^{\prime}}{h_{G+1}}+\frac{G+2}{r}v_{1}^{(1)(2,0)}-\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G+1}}u_{1}^{(1)(2,0)}+\sigma_{s}u_{1}^{(1)(1,0)}=0;\\ &v_{2}^{(1)(2,0)\prime}+v_{2}^{(1)(2,0)}\frac{h_{G-1}^{\prime}}{h_{G-1}}-\frac{G-1}{r}v_{2}^{(1)(2,0)}+\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G-1}}u_{2}^{(1)(2,0)}-\sigma_{s}u_{2}^{(1)(1,0)}=0;\\ &u_{1}^{(1)(2,0)\prime}+u_{1}^{(1)(2,0)}\frac{h_{G}^{\prime}}{h_{G}}-\frac{G}{r}u_{1}^{(1)(2,0)}+\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G+1}}{\mathrm{~h}_{G}}v_{1}^{(1)(2,0)}+\sigma_{s}v_{1}^{(1)(1,0)}=0;\\ &u_{2}^{(1)(2,0)\prime}+u_{2}^{(1)(2,0)}\frac{h_{G}^{\prime}}{h_{G}}+\frac{G+1}{r}u_{2}^{(1)(2,0)}-\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G-1}}{h_{G}}v_{2}^{(1)(2,0)}-\sigma_{s}v_{2}^{(1)(1,0)}=0;\end{aligned}\right$. (79)
    $\left\{\begin{aligned} &v_{1}^{(2)(2,0)\prime}+v_{1}^{(2)(2,0)}\frac{h_{G+1}^{\prime}}{h_{G+1}}+\frac{G+2}{r}v_{1}^{(2)(2,0)}-\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G+1}}u_{1}^{(2)(2,0)}+\sigma_{s}u_{1}^{(2)(1,0)}=0;\\ &v_{2}^{(2)(2,0)\prime}+v_{2}^{(2)(2,0)}\frac{h_{G-1}^{\prime}}{h_{G-1}}-\frac{G-1}{r}v_{2}^{(2)(2,0)}+\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G-1}}u_{2}^{(2)(2,0)}-\sigma_{s}u_{2}^{(2)(1,0)}=0;\\ &u_{1}^{(2)(2,0)\prime}+u_{1}^{(2)(2,0)}\frac{h_{G}^{\prime}}{h_{G}}-\frac{G}{r}u_{1}^{(2)(2,0)}+\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G+1}}{\mathrm{~h}_{G}}v_{1}^{(2)(2,0)}+\sigma_{s}v_{1}^{(2)(1,0)}=0;\\ &u_{2}^{(2)(2,0)\prime}+u_{2}^{(2)(2,0)}\frac{h_{G}^{\prime}}{h_{G}}+\frac{G+1}{r}u_{2}^{(2)(2,0)}-\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G-1}}{\mathrm{~h}_{G}}v_{2}^{(2)(2,0)}-\sigma_{s}v_{2}^{(2)(1,0)}=0;\end{aligned}\right$. (80)
@@
Appendix B Calculations of Born subtractions of phase shifts by the second order differential equations
@@
Appendix C Comparisons between results from Born subtraction and fake boson methods

### v4

@@ Appendix B Calculations of Born subtractions of phase shifts by the first order differential equations
    $\left\{\begin{aligned} &v_{1}^{(1)(1,0)\prime}+v_{1}^{(1)(1,0)}\frac{h_{G+1}^{\prime}}{h_{G+1}}+\frac{G+2}{r}v_{1}^{(1)(1,0)}-\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G+1}}u_{1}^{(1)(1,0)}+\beta\sigma_{s}=0;\\ &v_{2}^{(1)(1,0)\prime}+v_{2}^{(1)(1,0)}\frac{h_{G-1}^{\prime}}{h_{G-1}}-\frac{G-1}{r}v_{2}^{(1)(1,0)}+\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G-1}}u_{2}^{(1)(1,0)}=0\text{; }\\ &u_{1}^{(1)(1,0)\prime}+u_{1}^{(1)(1,0)}\frac{h_{G}^{\prime}}{h_{G}}-\frac{G}{r}u_{1}^{(1)(1,0)}+\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G+1}}{h_{G}}v_{1}^{(1)(1,0)}+\sigma_{s}=0;\\ &u_{2}^{(1)(1,0)\prime}+u_{2}^{(1)(1,0)}\frac{h_{G}^{\prime}}{h_{G}}+\frac{G+1}{r}u_{2}^{(1)(1,0)}-\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G-1}}{h_{G}}v_{2}^{(1)(1,0)}=0;\end{aligned}\right$. (94)
    $\left\{\begin{aligned} &v_{1}^{(2)(1,0)\prime}+v_{1}^{(2)(1,0)}\frac{h_{G+1}^{\prime}}{h_{G+1}}+\frac{G+2}{r}v_{1}^{(2)(1,0)}-\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G+1}}u_{1}^{(2)(1,0)}=0;\\ &v_{2}^{(2)(1,0)\prime}+v_{2}^{(2)(1,0)}\frac{h_{G-1}^{\prime}}{h_{G-1}}-\frac{G-1}{r}v_{2}^{(2)(1,0)}+\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G-1}}u_{2}^{(2)(1,0)}-\beta\sigma_{s}=0;\\ &u_{1}^{(2)(1,0)\prime}+u_{1}^{(2)(1,0)}\frac{h_{G}^{\prime}}{h_{G}}-\frac{G}{r}u_{1}^{(2)(1,0)}+\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G+1}}{h_{G}}v_{1}^{(2)(1,0)}=0;\\ &u_{2}^{(2)(1,0)\prime}+u_{2}^{(2)(1,0)}\frac{h_{G}^{\prime}}{h_{G}}+\frac{G+1}{r}u_{2}^{(2)(1,0)}-\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G-1}}{h_{G}}v_{2}^{(2)(1,0)}-\sigma_{s}=0;\end{aligned}\right$. (95)
@@
    $\left\{\begin{aligned} &v_{1}^{(1)(2,0)\prime}+v_{1}^{(1)(2,0)}\frac{h_{G+1}^{\prime}}{h_{G+1}}+\frac{G+2}{r}v_{1}^{(1)(2,0)}-\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G+1}}u_{1}^{(1)(2,0)}+\sigma_{s}u_{1}^{(1)(1,0)}=0;\\ &v_{2}^{(1)(2,0)\prime}+v_{2}^{(1)(2,0)}\frac{h_{G-1}^{\prime}}{h_{G-1}}-\frac{G-1}{r}v_{2}^{(1)(2,0)}+\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G-1}}u_{2}^{(1)(2,0)}-\sigma_{s}u_{2}^{(1)(1,0)}=0;\\ &u_{1}^{(1)(2,0)\prime}+u_{1}^{(1)(2,0)}\frac{h_{G}^{\prime}}{h_{G}}-\frac{G}{r}u_{1}^{(1)(2,0)}+\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G+1}}{h_{G}}v_{1}^{(1)(2,0)}+\sigma_{s}v_{1}^{(1)(1,0)}=0;\\ &u_{2}^{(1)(2,0)\prime}+u_{2}^{(1)(2,0)}\frac{h_{G}^{\prime}}{h_{G}}+\frac{G+1}{r}u_{2}^{(1)(2,0)}-\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G-1}}{h_{G}}v_{2}^{(1)(2,0)}-\sigma_{s}v_{2}^{(1)(1,0)}=0;\end{aligned}\right$. (96)
    $\left\{\begin{aligned} &v_{1}^{(2)(2,0)\prime}+v_{1}^{(2)(2,0)}\frac{h_{G+1}^{\prime}}{h_{G+1}}+\frac{G+2}{r}v_{1}^{(2)(2,0)}-\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G+1}}u_{1}^{(2)(2,0)}+\sigma_{s}u_{1}^{(2)(1,0)}=0;\\ &v_{2}^{(2)(2,0)\prime}+v_{2}^{(2)(2,0)}\frac{h_{G-1}^{\prime}}{h_{G-1}}-\frac{G-1}{r}v_{2}^{(2)(2,0)}+\left(\varepsilon-g\sigma_{v}\right)\frac{h_{G}}{h_{G-1}}u_{2}^{(2)(2,0)}-\sigma_{s}u_{2}^{(2)(1,0)}=0;\\ &u_{1}^{(2)(2,0)\prime}+u_{1}^{(2)(2,0)}\frac{h_{G}^{\prime}}{h_{G}}-\frac{G}{r}u_{1}^{(2)(2,0)}+\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G+1}}{h_{G}}v_{1}^{(2)(2,0)}+\sigma_{s}v_{1}^{(2)(1,0)}=0;\\ &u_{2}^{(2)(2,0)\prime}+u_{2}^{(2)(2,0)}\frac{h_{G}^{\prime}}{h_{G}}+\frac{G+1}{r}u_{2}^{(2)(2,0)}-\left(\varepsilon+g\sigma_{v}\right)\frac{h_{G-1}}{h_{G}}v_{2}^{(2)(2,0)}-\sigma_{s}v_{2}^{(2)(1,0)}=0;\end{aligned}\right$. (97)
@@
Appendix C Calculations of Born subtractions of phase shifts by the second order differential equations
@@
Appendix D Comparisons between results from Born subtraction and fake boson methods